OpenStudy (anonymous):

For what number of DVDs would the total cost of buying the DVDs without a club membership equal the total cost of joining the club and buying the DVDs with a club membership? (More information in a second)

OpenStudy (anonymous):

So this is an open response question, it's a little harder to ask for help, so... Info given- Each DVD costs $16 To buy a membership it's $40, and from then on the DVDs for members are only $12 a) What's the total cost of buying 5 DVDs at the store? $80 b) Write an equation that represents y, the total cost of buying x DVDs at the video store. y = 16x c) What's the total cost of joining the club and buying 5 DVDs? $100 d) Write an equation that represents y, the total cost to a club member, including the membership fee to join and buying x DVDs. y=40 + 12x

jimthompson5910 (jim_thompson5910):

ok where are you stuck?

OpenStudy (anonymous):

This one For what number of DVDs would the total cost of buying the DVDs without a club membership equal the total cost of joining the club and buying the DVDs with a club membership?

OpenStudy (anonymous):

Would I do the 16x = 40 + 12x thing?

jimthompson5910 (jim_thompson5910):

The total cost to buy x DVDs with no membership is y = 16x The total cost to buy x DVDs with membership is y = 12x+40 so this means... y = 12x + 40 ... start with second equation 16x = 12x + 40 ... plug in y = 16x now solve for x

jimthompson5910 (jim_thompson5910):

yes exactly

jimthompson5910 (jim_thompson5910):

you're allowed to this because you want to know how many DVDs (x) you can buy for both to make the cost (y) equal, so basically x and y are the same in each equation in this case

OpenStudy (anonymous):

16x = 12x + 40 4x = 40 x = 10?

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

so you're done optionally you can find the equal cost, but they just want the number of DVDs

OpenStudy (anonymous):

Thank you!

jimthompson5910 (jim_thompson5910):

you're welcome